If a vendor charges p dollars each for rugby shirts then he expects to sell
2000 - 100p shirts at a tournament.
a)Find a polynomial R(p) that represents the total revenue when the shirts are p dollars.
b) Find R(5), R(10), and R(20)
I don't understand what they are asking and how to set up the polynomial....please HELP!!!

Answer by robertb(5830)

a) R(p) = p(2000 - 100p), because revenue = (unit price)*(total number of sales).

b) R(5) = 5*(2000 - 500) =5*1500 = 7,500 $.
R(10) = 10*(2000 - 1000) = 10*1000 = 10,000 $
R(20) = 20(2000 - 2000) = 0$ (This means that if he charges a unit price of $20, the vendor won't sell anything.)